# Integral points and volume of integral-integral affine manifolds

### v2

@@ 1. Introduction
On an $n$-dimensional manifold $M$, an atlas whose transition maps locally have the form $x\mapsto Ax+b$ with $A\in\mathsf{GL}(n,\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$ defines an integral-integral affine structure [8] (also called a strongly integral affine structure [14]), which determines a measure on $M$ and a subset $M_{\mathbb{Z}}\subset M$ of integral points. (See Definition 2.5, Remark 2.6, Remark 2.7, and Definition 2.12.)
@@
Theorem 1.1 appeared as the “Downstairs Theorem” in [8]. Its proof was surprizingly nontrivial, and involved intersection theory on a bundle of tori over MM; [^1] In [8], the integral-integral manifold was called $B$; the letter $M$ was used for various types of torus bundles over $B$. (A rough sketch for this proof already existed in the literature, but the details turned out to require a subtle cohomological argument, which was provided in [8].)
In this paper we provide a more elementary proof of Theorem 1.1 that relies on rational Ehrhart theory in $\mathbb{R}^{n}$ [4, 1]. Here is a sketch of our argument. We introduce the lattice count function, which associates to each $m\in\mathbb{N}$ the number $L_{M}(m)$ of points in $M$ whose coordinates are in $m^{-1}\mathbb{Z}$. By “decomposing” $M$ into rational polytopes contained in charts,222 We do not give a polyhedral decomposition of $M$. Instead, we cover $M$ by overlapping rational polytopes, and we apply an inclusion-exclusion argument. we conclude that $L_{M}(m)$ is a quasi-polynomial in $m$. By “decomposing” $M$ using a smooth partition of unity supported in charts, we conclude that $L_{M}(m)-\text{vol}(M)m^{n}=o(1)$ as $m\to\infty$. From these two results, we conclude that $L_{M}(m)=\text{vol}(M)m^{n}$ on the nose, yielding Theorem 1.1 when $m=1$.
We note that we do not know if the universal cover of a compact integral-integral affine manifold must be isomorphic to $\mathbb{R}^{n}$ with its standard integral-integral-affine structure. This is a special case of the Markus conjecture [11] (see the discussion in [5]).
@@ Acknowledgement
Y. Karshon’s research is partly funded by the Natural Sciences and Engineering Research Council of Canada Discovery RGPIN-2024-05798 and by the United States – Israel Binational Science Foundation Grant 2021730.
@@ 2. Preliminaries
2.2 Definition.
The set of integral-integral affine maps of $\mathbb{R}^{n}$ is
$\text{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})\coloneq\{f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},\;f(x)=b+Ax\mid b\in\mathbb{Z}^{n},\;A\in\mathsf{GL}_{n}(\mathbb{Z})\}\,.$
@@
for M is called an integral-integral affine atlas if $\varphi_{j}\circ\varphi_{i}^{-1}$ is locally integral-integral affine on $\varphi_{i}(U_{i}\cap U_{j})$ for every $i,j\in I$. The integral-integral affine atlases are ordered by inclusion. An integral-integral affine structure on a manifold $M$ is a maximal integral-integral affine atlas. An integral integral affine manifold is a manifold equipped with an integral-integral affine structure. An integral chart is an element of the maximal integral-integral affine atlas of $M$.
@@
The definition of a polytope in [6] is equivalent to our Definition 2.8, by [6, Theorem 1.26]. The result is then in [6, Proposition 1.69]. ∎
@@ 3. Main Theorems
Because every open cover of $M$ has a refinement that is a good open cover of $M$ [3, Propositions 5.1 and 5.2], there exists a good open cover $\mathcal{W}$} of $M$ such that every element of $\mathcal{W}$} is contained in the domain of an integral chart.
@@
Because $f$ is a compactly supported smooth function, it is a Schwartz function. Because Fourier transform maps Schwartz functions to Schwartz functions [15, p.181, Corollary 2.2], the Fourier transform
@@
By the Poisson summation formula [13, p.88, Theorem 3.44],
@@
Asymptotic expansions of a similar nature appeared in [7], [10] and [2]. The idea to take advantage of quasi-polynomial behavior to show that an asymptotic result is exact appeared in [12]. In the future, it might be interesting to seek generalizations to integral-integral affine manifolds of Euler-Maclaurin formulas relating sums to integrals [9, 2].

### v3

@@ 1. Introduction
On an $n$-dimensional manifold $M$, an atlas whose transition maps locally have the form
(1.1) $x\mapsto Ax+b\quad\text{ with $A\in\mathsf{GL}(n,\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$ }$
defines an integral-integral affine structure [undefi] (also called a strongly integral affine structure [undefo]). This structure determines a measure on $M$ and a subset $M_{\mathbb{Z}}\subset M$ of integral points. (See Definition 2.5, Remark 2.6, Remark 2.7, and Definition 2.12.)
For example, if $\Gamma$ is group of transformations of $\mathbb{R}^{n}$ of the form (1.1) whose action on $\mathbb{R}^{n}$ is free and proper then the quotient $M:=\mathbb{R}^{n}/\Gamma$ inherits an integral-integral affine structure. We do not know if every compact integral-integral affine manifold must be isomorphic to such a quotient; this is a special case of the Markus conjecture [undefl] (see the discussion in [undefe]).
In this paper we give an elementary proof of the following theorem:
@@
Our motivation for this project comes from geometric quantization. Roughly, geometric quantization provides recipes to pass from a symplectic manifold $(X,\omega)$, which serves as a model for phase space in classical mechanics, to a Hilbert space $\mathcal{H}$, whose elements we view as wave functions in the corresponding quantum mechanical system. Such recipes involve the choice of a polarization on $X$. Properties of the quantum mechanical system often turn out to be independent of this choice; this phenomenon is called “independence of polarization”.
One example of a polarization is a compatible Kähler structure on $X$; some recipes allow a compatible almost complex structure; the dimension of the corresponding Hilbert space is typically given by the Riemann-Roch number. Another example of a polarization is a Lagrangian fibration $X\to B$. In the presence of prequantization, the base $B$ then acquires an integral-integral affine structure, and the dimension of the corresponding Hilbert space that arises from this polarization is given by the number of integral points. The Riemann-Roch number of $(X,\omega)$ then coincides with the volume of $B$; thus, Theorem 1.2 becomes an instance of “independence of polarization”.
Theorem 1.2 appeared as the “Downstairs Theorem” in [undefi]. Its proof was surprizingly nontrivial, and involved intersection theory on a bundle of tori over $M$. (Warning: in [undefi], the integral-integral manifold was called $B$; the letter $M$ was used for various types of torus bundles over $B$.) The proof in [undefi] was inspired by ideas that were sketched inside papers on geometric quantization and on mirror symmetry ([undef, Proof of Thm. 4.1]; [undeff, Conjecture 4.9]); the details turned out to require a subtle cohomological argument, which was provided in [undefi]. In light of this earlier work, the elementary proof that we provide in this current paper is particularly satisfying.
In this paper we provide a more elementary proof of Theorem 1.2 that relies on rational Ehrhart theory in $\mathbb{R}^{n}$ [undefd, undefa]. Here is a sketch of our argument. We introduce the lattice count function, which associates to each $m\in\mathbb{N}$ the number $L_{M}(m)$ of points in $M$ whose coordinates are in $m^{-1}\mathbb{Z}$. For a rational convex polytope $P$ in $\mathbb{R}^{n}$, the number of points in $P\cap m^{-1}\mathbb{Z}^{n}$ is quasi-polynomial in $m$ (Definition 2.10), by rational Ehrhart theory. For a compact integral-integral affine manifold $M$, we do not know how to decompose $M$ into rational convex polytopes; however, we can cover $M$ by overlapping rational convex polytopes, and apply an inclusion-exclusion argument, to conclude that $L_{M}(m)$ is a quasi-polynomial in $m$. On the other hand, by “decomposing” $M$ using a smooth partition of unity supported in charts, we conclude that $L_{M}(m)-\text{vol}(M)m^{n}=o(1)$ as $m\to\infty$. From these two results, we conclude that $L_{M}(m)=\text{vol}(M)m^{n}$ on the nose, yielding Theorem 1.2.
@@ Acknowledgement
Y. Karshon’s research is partly funded by the Natural Sciences and Engineering Research Council of Canada Discovery RGPIN-2024-05798 and by the United States – Israel Binational Science Foundation Grant 2021730. O. Elisha is partly supported by Yaron Ostrover’s ISF Grant 938/22.
@@ 2. Preliminaries
2.2 Notation.
We denote by $\textbf{Aff}_{\mathbb{Z}}(\mathbb{R}^{n})$ the set of maps $\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $x\mapsto Ax+b$ with $A\in\mathsf{GL}(n,\mathbb{Z})$ and $b\in\mathbb{Z}^{n}$.
@@
for M is called an integral-integral affine atlas if $\varphi_{j}\circ\varphi_{i}^{-1}$ is locally integral-integral affine on $\varphi_{i}(U_{i}\cap U_{j})$ for every $i,j\in I$. The integral-integral affine atlases are ordered by inclusion. An integral-integral affine structure on a manifold $M$ is a maximal integral-integral affine atlas. An integral-integral affine manifold is a manifold equipped with an integral-integral affine structure. An integral chart is an element of the maximal integral-integral affine atlas of $M$.
@@
The definition of a polytope in [undefg] is equivalent to our Definition 2.8, by [undefg, Theorem 1.26]. The result is then in [undefg, Proposition 1.69]. ∎
@@ 3. Main Theorems
Because every open cover of $M$ has a refinement that is a good open cover of $M$ [undefc, Propositions 5.1 and 5.2], there exists a good open cover $\mathcal{W}$} of $M$ such that every element of $\mathcal{W}$} is contained in the domain of an integral chart.
@@
Because $f$ is a compactly supported smooth function, it is a Schwartz function. Because Fourier transform maps Schwartz functions to Schwartz functions [undefp, p.181, Corollary 2.2], the Fourier transform
@@
By the Poisson summation formula [undefn, p.88, Theorem 3.44],
@@
Asymptotic expansions of a similar nature appeared in [undefh], [undefk] and [undefb]. The idea to take advantage of quasi-polynomial behavior to show that an asymptotic result is exact appeared in [undefm]. In the future, it might be interesting to seek generalizations to integral-integral affine manifolds of Euler-Maclaurin formulas relating sums to integrals [undefj, undefb].